OpenStudy (anonymous):

solve over [0 degrees,360 degrees): tan2x-sqrt3=0

OpenStudy (anonymous):

is that that tan(2x) or tan(x) ^2 ?

OpenStudy (anonymous):

tan(2x) sorry

OpenStudy (anonymous):

tan(2x) = sqrt3 2x = tan-1(sqrt3) 2x = pi/3

OpenStudy (anonymous):

x = pi/6

OpenStudy (anonymous):

which is 30 degrees

OpenStudy (anonymous):

so, if x = pi/6, that means that the two angles would be 30 degrees, and 330 degrees, right?

OpenStudy (anonymous):

over (0 to 360 degrees), right?

OpenStudy (anonymous):

Well also, since tan funtion repeats every 180 degrees. Then 30+180=210* is also a valid answer

OpenStudy (anonymous):

 lol epic sketch